# Adam Exploits $\ell_{\infty}$-geometry of Loss Landscape via Coordinate-wise Adaptivity

### v1

@@ C.1 Proof for convergence rate of blockwise $\mathtt{Adam}$
$\mathbb{E}\max_{b\in[B]}v_{T,b}+\epsilon$ $\leq\epsilon+C+4(1-\beta_{2})\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\sum_{s=1}^{t}\beta_{2}^{t-s}G$
$\leq\epsilon+C+4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}G=G^{\prime}$
$\leq 2\epsilon+2C+8\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T^{2}+32(v_{0}+\epsilon)\left(\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T\frac{\beta_{2}}{(1-\beta_{2})(v_{0}+\epsilon)}\right)^{2}$
@@
$\leq$ $\ln{\frac{2\epsilon+2C+8\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T^{2}+32(v_{0}+\epsilon)\left(\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T\frac{\beta_{2}}{(1-\beta_{2})(v_{0}+\epsilon)}\right)^{2}}{v_{0}+\epsilon}}$
$\leq$ $\ln{\left[2\left(1+\frac{\sum_{b=1}^{B}\sigma_{b}^{2}+2\left\|\nabla L({\bm{x}}_{0})\right\|_{\Phi}^{2}}{v_{0}+\epsilon}+\frac{2\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T^{2}}{v_{0}+\epsilon}+\frac{4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T\beta_{2}}{(1-\beta_{2})(v_{0}+\epsilon)}\right)^{2}\right]}$
$\leq$ $2\ln\left(1+\frac{\sum_{b=1}^{B}\sigma_{b}^{2}+\left\|\nabla L({\bm{x}}_{0})\right\|_{\Phi}^{2}+\max_{b\in[B]}H_{b}^{2}\eta^{2}T(T+\frac{1}{1-\beta_{2}})}{v_{0}+\epsilon}\right)+\ln{32}$
@@
In a single step, we can apply Lemma 3.12 and have that
$L({\bm{x}}_{t})-L({\bm{x}}_{t-1})$ $\leq\nabla L({\bm{x}}_{t-1})^{\top}({\bm{x}}_{t}-{\bm{x}}_{t-1})+\frac{1}{2}\sum_{b=1}^{B}\frac{H_{b}}{d_{b}}\sum_{\Phi(i)=b}\left(x_{t,i}-x_{t-1,i}\right)^{2}$
$=-\eta\sum_{i=1}^{d}\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{v_{t,\Phi(i)}+\epsilon}}+\frac{1}{2}\eta^{2}\sum_{b=1}^{B}\frac{H_{b}}{d_{b}}\frac{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}}{v_{t,b}+\epsilon}.$
@@
$\mathbb{E}\left[L({\bm{x}}_{T})-L({\bm{x}}_{0})\right]$ $\leq-\mathbb{E}\left[\eta\sum_{i=1}^{d}\sum_{t=1}^{T}\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{v_{t,\Phi(i)}+\epsilon}}\right]+\frac{1}{2}\eta^{2}\mathbb{E}\left[\sum_{b=1}^{B}\frac{H_{b}}{d_{b}}\sum_{t=1}^{T}\frac{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}}{v_{t,b}+\epsilon}\right]$
$\leq-\mathbb{E}\left[\eta\sum_{i=1}^{d}\sum_{t=1}^{T}\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{v_{t,\Phi(i)}+\epsilon}}\right]+\frac{1}{2}\eta^{2}\mathbb{E}\left[\sum_{b=1}^{B}H_{b}\left(T+\frac{\beta_{2}}{1-\beta_{2}}\ln{\frac{v_{T,b}+\epsilon}{v_{0,b}+\epsilon}}\right)\right].$
@@
$\frac{1}{T}\mathbb{E}\left[\sum_{i=1}^{d}\sum_{t=1}^{T}\frac{\bar{g}_{t,i}^{2}}{\sqrt{\tilde{v}_{t,\Phi(i)}+\epsilon}}\right]$ $\leq\frac{2}{\eta T}\mathbb{E}\left[L({\bm{x}}_{0})-L({\bm{x}}_{T})\right]+\frac{\eta}{T}\mathbb{E}\left[\sum_{b=1}^{B}H_{b}\left(T+\frac{\beta_{2}}{1-\beta_{2}}\ln{\frac{v_{T,b}+\epsilon}{v_{0,b}+\epsilon}}\right)\right]$
@@
$\leq\frac{2}{\eta T}\mathbb{E}\left[L({\bm{x}}_{0})-L({\bm{x}}_{T})\right]+\eta\sum_{b=1}^{B}H_{b}+\sqrt{1-\beta_{2}}\sum_{b=1}^{B}d_{b}\sigma_{b}$
$+\frac{\beta_{2}}{T(1-\beta_{2})}\left(\eta\sum_{b=1}^{B}H_{b}+\sqrt{1-\beta_{2}}\sum_{b=1}^{B}\sigma_{b}\right)\max_{b\in[B]}\mathbb{E}\ln{\frac{v_{T,b}+\epsilon}{v_{0,b}+\epsilon}}$
$\leq\frac{2}{\eta T}\mathbb{E}\left[L({\bm{x}}_{0})-L({\bm{x}}_{T})\right]+\eta\sum_{b=1}^{B}H_{b}+\sqrt{1-\beta_{2}}\sum_{b=1}^{B}d_{b}\sigma_{b}$
$+\frac{\beta_{2}}{T(1-\beta_{2})}\left(\eta\sum_{b=1}^{B}H_{b}+\sqrt{1-\beta_{2}}\sum_{b=1}^{B}d_{b}\sigma_{b}\right)\ln{\frac{\mathbb{E}\max_{b\in[B]}v_{T,b}+\epsilon}{v_{0}+\epsilon}}$
@@
$E$ $=\frac{2}{\eta T}\mathbb{E}\left[L({\bm{x}}_{0})-L({\bm{x}}_{T})\right]+\eta\sum_{b=1}^{B}H_{b}+\sqrt{1-\beta_{2}}\sum_{b=1}^{B}d_{b}\sigma_{b}+\frac{\beta_{2}}{T(1-\beta_{2})}\left(\eta\sum_{b=1}^{B}H_{b}+\sqrt{1-\beta_{2}}\sum_{b=1}^{B}d_{b}\sigma_{b}\right)F,$
@@
$F=2\ln\left(1+\frac{\sum_{b=1}^{B}\sigma_{b}^{2}+\left\|\nabla L({\bm{x}}_{0})\right\|_{\Phi}^{2}+\max_{b\in[B]}H_{b}^{2}\eta^{2}T(T+\frac{1}{1-\beta_{2}})}{v_{0}+\epsilon}\right)+\ln{32}.$
@@ D.2 Computation of matrix norms
It is computationally infeasible to get the full Hessian matrix and directly compute norms of it. Instead we leverage Hessian vector product function in Jax to probe the Hessian matrix. We use Lanczos algorithm to estimate spectral norm. The estimation of $(1,1)$-norm relies on the properties of Cauchy distribution. Given $a=[a_{1},\cdots,a_{n}]$ and i.i.d. standard Cauchy variables $X_{1},\cdots,X_{n}$, $\sum_{i=1}^{n}a_{i}X_{i}$ is Cauchy distributed with mean 00 and scale $\sum_{i=1}^{n}\left|a_{i}\right|$. For a single value Cauchy distribution with mean 00 and scaling $\gamma$, an estimator for $\gamma$ is the median of the absolute values of all the samples.
Therefore, we propose Algorithm 4 to estimate the sum of absolute values for each row and sum over all the rows to get $(1,1)$-norm of Hessian matrix. We choose $n=50$ for all the measurement experiments. We also prove a concentration inequality in Theorem D.1.
Algorithm 4 Estimation of $(1,1)$-Norm of Hessian, $\nabla^{2}L(\theta)$
1:Number of Cauchy vectors $n$, parameter $\theta\in\mathbb{R}^{d}$, loss $L$
2:for $i=1\to n$ :
@@
4:   $\mathbf{H}_{:,i}\leftarrow\nabla^{2}L(\theta)\cdot v^{(i)}$ (Using hessian-vector product)
5:return $\sum_{j=1}^{d}\mathrm{median}(\mathrm{abs}(\mathbf{H}_{j,:}))$
Theorem D.1.
For the estimate in Algorithm 4, it holds that
$P\left(\left|\sum_{j=1}^{d}\textup{median}(\left|\mathbf{H}_{j,:}\right|)-\left\|\nabla^{2}L(\theta)\right\|_{1,1}\right|\geq\epsilon\left\|\nabla^{2}L(\theta)\right\|_{1,1}\right)\leq 2d\cdot\exp{\left(-\frac{8n}{25\pi^{2}}\epsilon^{2}\right)}$
for every $\epsilon\in(0,1)$.
Proof of Theorem D.1.
[…]
$P(\left|M_{n}-1\right|\geq\epsilon)\leq 2\exp{\left(-\frac{8n}{25\pi^{2}}\epsilon^{2}\right)}.$ (11)
Given $X\sim\textrm{Cauchy}[0,1]$, the cumulative distribution function of $\left|X\right|$ is $F(x)=\frac{2}{\pi}\textrm{arctan}(x)$ and the median of $\left|X\right|$ is $1$ because $F(1)=0.5$. For a fixed $\epsilon\in(0,1)$, define $p_{1}=\frac{2}{\pi}\textrm{arctan}(1-\epsilon)$ to be the probability that $\left|X\right|$ is smaller than $1-\epsilon$ and $S=\sum_{i=1}^{n}\mathbf{1}_{\left|X_{i}\right|\leq 1-\epsilon}$. Since $\mathbf{1}_{\left|X_{i}\right|\leq 1-\epsilon}$ follows i.i.d. Bernoulli distribution with $p_{1}$, $S\sim\textrm{Bin}(n,p_{1})$.
$M_{n}\leq 1-\epsilon$ if and only if at least $\frac{n+1}{2}$ $X_{i}$’s are smaller than $1-\epsilon$. And we can apply Hoeffding’s inequality on $S$ and get that
$P(M_{n}\leq 1-\epsilon)$ $=P(S\geq\frac{n+1}{2})\leq P(S\geq\frac{n}{2})$
@@
$=\exp{\left(-2n\left(\frac{1}{2}-\frac{2}{\pi}\textrm{arctan}(1-\epsilon)\right)^{2}\right)}.$
With similar derivation, we can also get that
$P(M_{n}\geq 1+\epsilon)\leq\exp{\left(-2n\left(\frac{1}{2}-\frac{2}{\pi}\textrm{arctan}(1+\epsilon)\right)^{2}\right)}.$
When $\epsilon\in[-1,1]$, we have that
$\left(\frac{1}{2}-\frac{2}{\pi}\textrm{arctan}(1+\epsilon)\right)^{2}=\frac{4}{\pi^{2}}\left(\textrm{arctan}(1)-\textrm{arctan}(1+\epsilon)\right)^{2}\geq\frac{4}{25\pi^{2}}\epsilon^{2}.$
because $\textrm{arctan}^{\prime}(x)=\frac{1}{1+x^{2}}$ is always greater than $\frac{1}{5}$ for $x\in[0,2]$. Then we can have Equation 11.
@@
$P\left(\left|\textrm{median}(\left|H_{j,1}\right|,\cdots,\left|H_{j,n}\right|)-\sum_{k=1}^{d}\left|\nabla^{2}L(\theta)_{j,k}\right|\right|\geq\epsilon\sum_{k=1}^{d}\left|\nabla^{2}L(\theta)_{j,k}\right|\right)\leq 2\exp{\left(-\frac{8n}{25\pi^{2}}\epsilon^{2}\right)}.$
And we can sum the tail probability over all the rows $j$ and get that
$P\left(\left|\sum_{j=1}^{d}\textrm{median}(\left|\mathbf{H}_{j,:}\right|)-\left\|\nabla^{2}L(\theta)\right\|_{1,1}\right|\geq\epsilon\left\|\nabla^{2}L(\theta)\right\|_{1,1}\right)\leq 2d\exp{\left(-\frac{8n}{25\pi^{2}}\epsilon^{2}\right)}.$

### v3

@@ Appendix C Proof for Convergence Rate of Blockwise $\mathtt{Adam}$
$G\leq 2\epsilon+2C+8\sum_{b\in[B]}H_{b}^{2}d_{b}\eta^{2}T^{2}+32(v_{0}+\epsilon)\left(\sum_{b\in[B]}H_{b}^{2}d_{b}\eta^{2}T\frac{\beta_{2}}{(1-\beta_{2})(v_{0}+\epsilon)}\right)^{2}$
@@
$\leq$ $\ln{\frac{2\epsilon+2C+8\sum_{b\in[B]}H_{b}^{2}d_{b}\eta^{2}T^{2}+32(v_{0}+\epsilon)\left(\sum_{b\in[B]}H_{b}^{2}d_{b}\eta^{2}T\frac{\beta_{2}}{(1-\beta_{2})(v_{0}+\epsilon)}\right)^{2}}{v_{0}+\epsilon}}$
$\leq$ $\ln{\left[2\left(1+\frac{\sum_{b=1}^{B}\sigma_{b}^{2}+2\left\|\nabla L({\bm{x}}_{0})\right\|_{\Phi}^{2}}{v_{0}+\epsilon}+\frac{2\sum_{b\in[B]}H_{b}^{2}d_{b}\eta^{2}T^{2}}{v_{0}+\epsilon}+\frac{4\sum_{b\in[B]}H_{b}^{2}d_{b}\eta^{2}T\beta_{2}}{(1-\beta_{2})(v_{0}+\epsilon)}\right)^{2}\right]}$
$\leq$ $2\ln\left(1+\frac{\sum_{b=1}^{B}\sigma_{b}^{2}+\left\|\nabla L({\bm{x}}_{0})\right\|_{\Phi}^{2}+\sum_{b\in[B]}H_{b}^{2}d_{b}\eta^{2}T(T+\frac{1}{1-\beta_{2}})}{v_{0}+\epsilon}\right)+\ln{32}$
@@
From Definition 3.9, there exists a diagonal matrix ${\bm{H}}$ that follows $\Phi$ and always dominates $\nabla^{2}L({\bm{x}})$ satisfying $H(L,\Phi)=\operatorname{Tr}({\bm{H}})=\sum_{b\in[B]}H_{b}d_{b}$. In a single step, we can have that
$L({\bm{x}}_{t})-L({\bm{x}}_{t-1})$ $\leq\nabla L({\bm{x}}_{t-1})^{\top}({\bm{x}}_{t}-{\bm{x}}_{t-1})+\frac{1}{2}\sum_{b=1}^{B}H_{b}\sum_{\Phi(i)=b}\left(x_{t,i}-x_{t-1,i}\right)^{2}$
$=-\eta\sum_{i=1}^{d}\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{v_{t,\Phi(i)}+\epsilon}}+\frac{1}{2}\eta^{2}\sum_{b=1}^{B}H_{b}\frac{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}}{v_{t,b}+\epsilon}.$
@@
$\mathbb{E}\left[L({\bm{x}}_{T})-L({\bm{x}}_{0})\right]$ $\leq-\mathbb{E}\left[\eta\sum_{i=1}^{d}\sum_{t=1}^{T}\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{v_{t,\Phi(i)}+\epsilon}}\right]+\frac{1}{2}\eta^{2}\mathbb{E}\left[\sum_{b=1}^{B}H_{b}\sum_{t=1}^{T}\frac{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}}{v_{t,b}+\epsilon}\right]$
$\leq-\mathbb{E}\left[\eta\sum_{i=1}^{d}\sum_{t=1}^{T}\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{v_{t,\Phi(i)}+\epsilon}}\right]+\frac{1}{2}\eta^{2}\mathbb{E}\left[\sum_{b=1}^{B}H_{b}d_{b}\left(T+\frac{\beta_{2}}{1-\beta_{2}}\ln{\frac{v_{T,b}+\epsilon}{v_{0,b}+\epsilon}}\right)\right].$
@@
$\frac{1}{T}\mathbb{E}\left[\sum_{i=1}^{d}\sum_{t=1}^{T}\frac{\bar{g}_{t,i}^{2}}{\sqrt{\tilde{v}_{t,\Phi(i)}+\epsilon}}\right]$ $\leq\frac{2}{\eta T}\mathbb{E}\left[L({\bm{x}}_{0})-L({\bm{x}}_{T})\right]+\frac{\eta}{T}\mathbb{E}\left[\sum_{b=1}^{B}H_{b}d_{b}\left(T+\frac{\beta_{2}}{1-\beta_{2}}\ln{\frac{v_{T,b}+\epsilon}{v_{0,b}+\epsilon}}\right)\right]$
@@
$\leq\frac{2}{\eta T}\mathbb{E}\left[L({\bm{x}}_{0})-L({\bm{x}}_{T})\right]+\eta\sum_{b=1}^{B}H_{b}d_{b}+\sqrt{1-\beta_{2}}\sum_{b=1}^{B}d_{b}\sigma_{b}$
$+\frac{\beta_{2}}{T(1-\beta_{2})}\left(\eta\sum_{b=1}^{B}H_{b}d_{b}+\sqrt{1-\beta_{2}}\sum_{b=1}^{B}\sigma_{b}\right)\max_{b\in[B]}\mathbb{E}\ln{\frac{v_{T,b}+\epsilon}{v_{0,b}+\epsilon}}$
$\leq\frac{2}{\eta T}\mathbb{E}\left[L({\bm{x}}_{0})-L({\bm{x}}_{T})\right]+\eta\sum_{b=1}^{B}H_{b}d_{b}+\sqrt{1-\beta_{2}}\sum_{b=1}^{B}d_{b}\sigma_{b}$
$+\frac{\beta_{2}}{T(1-\beta_{2})}\left(\eta\sum_{b=1}^{B}H_{b}d_{b}+\sqrt{1-\beta_{2}}\sum_{b=1}^{B}d_{b}\sigma_{b}\right)\ln{\frac{\mathbb{E}\max_{b\in[B]}v_{T,b}+\epsilon}{v_{0}+\epsilon}}$
@@
$E$ $=\frac{2}{\eta T}\mathbb{E}\left[L({\bm{x}}_{0})-L({\bm{x}}_{T})\right]+\eta\sum_{b=1}^{B}H_{b}d_{b}+\sqrt{1-\beta_{2}}\sum_{b=1}^{B}d_{b}\sigma_{b}+\frac{\beta_{2}}{T(1-\beta_{2})}\left(\eta\sum_{b=1}^{B}H_{b}d_{b}+\sqrt{1-\beta_{2}}\sum_{b=1}^{B}d_{b}\sigma_{b}\right)F,$
@@
$F=2\ln\left(1+\frac{\sum_{b=1}^{B}\sigma_{b}^{2}+\left\|\nabla L({\bm{x}}_{0})\right\|_{\Phi}^{2}+\sum_{b\in[B]}H_{b}^{2}d_{b}\eta^{2}T(T+\frac{1}{1-\beta_{2}})}{v_{0}+\epsilon}\right)+\ln{32}.$
@@ D.2 Computation of matrix norms
As mentioned in Section 3.3, it is computationally infeasible to get the full Hessian matrix and directly compute norms of it. Instead we leverage Hessian vector product function in Jax to probe the Hessian matrix. We use Lanczos algorithm to estimate spectral norm.
We propose Algorithm 4 to estimate the sum of absolute values for each row and sum over all the rows to get $(1,1)$-norm of Hessian matrix. We first subsample a fixed batch of training data for estimating the $(1,1)$-norm of Hessian matrix. The high-level idea is to compute the matrix vector products between Hessian of training loss on this batch and a sequence of random Cauchy vectors. Then we take the $\ell_{1}$ norm of the coordinate-wise median of the resulting sequence of Hessian vector products. Because the Cauchy distribution is 1-stable, the resulting product is also a vector of Cauchy random variables, and the magnitude of each element equals to $\ell_{1}$ norm of the corresponding row of the Hessian. Thus with infinitely many samples, the $\ell_{1}$ norm of the coordinate-wise median converges almost surely to the $(1,1)$-norm of the Hessian.
We choose $n=200$ for the measurement experiments on GPT-2 and $n=50$ for the measurement experiments on ResNet18. We also prove a non-asymptotic high-probability multiplicative bound for the estimation error which depends mildly on the dimension $d$ in Theorem 4.1.
Algorithm 4 Estimation of $(1,1)$-Norm of Hessian, $\nabla^{2}L({\bm{x}})$
1:Number of Cauchy vectors $n$, parameter ${\bm{x}}\in\mathbb{R}^{d}$, loss $L$
2:for $i=1$ to $n$ :
@@
4:   $\mathbf{H}_{:,i}\leftarrow\nabla^{2}L({\bm{x}})\cdot v^{(i)}$ (Using hessian-vector product)
5:return $\sum_{j=1}^{d}\mathrm{median}(\left|\mathbf{H}_{j,:}\right|)$
First we prove that median of random variables following uniform distribution is sub-Gaussian.
Lemma D.1.
Suppose $Z_{1},\cdots,Z_{n}\overset{\mathrm{iid}}{\sim}\textup{Unif}([0,1])$. Then $P(\left|\textup{median}(Z_{1},\cdots,Z_{n})-\frac{1}{2}\right|\geq\epsilon)\leq 2\exp{\left(-2n\epsilon^{2}\right)}$ for any $\epsilon\geq 0$.
Proof of Lemma D.1.
Define $S=\sum_{i=1}^{n}\mathbf{1}_{Z_{i}\leq\frac{1}{2}-\epsilon}$. Since $\mathbf{1}_{Z_{i}\leq\frac{1}{2}-\epsilon}$ follows i.i.d. Bernoulli distribution with $p_{1}=\frac{1}{2}-\epsilon$, $S\sim\textup{Bin}(n,p_{1})$.
$M_{n}=\textup{median}(Z_{1},\cdots,Z_{n})\leq\frac{1}{2}-\epsilon$ if and only if at least $\frac{n+1}{2}$ $Z_{i}$’s are smaller than $\frac{1}{2}-\epsilon$. And we can apply Hoeffding’s inequality on $S$ and get that
$P(M_{n}\leq\frac{1}{2}-\epsilon)$ $=P(S\geq\frac{n+1}{2})\leq P(S\geq\frac{n}{2})$
@@
$=\exp{\left(-2n\epsilon^{2}\right)}.$
We can know $P(M_{n}\geq\frac{1}{2}+\epsilon)\leq\exp(-2n\epsilon^{2})$ from the symmetry of distribution. ∎
See 4.1
Proof of Theorem 4.1.
@@
$P\left(\left|\sum_{j=1}^{d}a_{j}\textup{median}(\left|Y_{j,1}\right|,\cdots,\left|Y_{j,n}\right|)-\sum_{j=1}^{d}a_{j}\right|\geq\epsilon\sum_{j=1}^{d}a_{j}\right)\leq 2d\exp(-\frac{n\Delta^{2}}{2})+2\exp{\left(-\frac{2n\epsilon^{2}\cos^{4}((1+\Delta)\frac{\pi}{4})}{\pi^{2}}\right)}.$ (12)
[…]
$P\left(\left|\sum_{j=1}^{d}a_{j}\textup{median}(\tan(X_{j,1}),\cdots,\tan(X_{j,n}))-\sum_{j=1}^{d}a_{j}\right|\geq\epsilon\sum_{j=1}^{d}a_{j}\right)\leq 2d\exp(-\frac{n\Delta^{2}}{2})+2\exp{\left(-\frac{2n\epsilon^{2}\cos^{4}((1+\Delta)\frac{\pi}{4})}{\pi^{2}}\right)}.$ (13)
for any $\{X_{j,k}\}$ such that $X_{j,1},\cdots,X_{j,n}\overset{\mathrm{iid}}{\sim}\textup{Unif}(0,\frac{\pi}{2})$ for any $j\in[d]$.
Fix $\Delta\in(0,1)$. We define
$f(x)=\begin{cases}\tan(x)&\text{if }\left|x-\frac{\pi}{4}\right|\leq\Delta\frac{\pi}{4},\\ \frac{1}{\cos^{2}((1-\Delta)\frac{\pi}{4})}(x-(1-\Delta)\frac{\pi}{4})+\tan((1-\Delta)\frac{\pi}{4})&\text{if }0<x<(1-\Delta)\frac{\pi}{4},\\ \frac{1}{\cos^{2}((1+\Delta)\frac{\pi}{4})}(x-(1+\Delta)\frac{\pi}{4})+\tan((1+\Delta)\frac{\pi}{4})&\text{if }(1+\Delta)\frac{\pi}{4}<x<\frac{\pi}{2}.\end{cases}$
Then $f(x)$ is a differentiable function on $(0,\frac{\pi}{2})$ that equals to $\tan(x)$ in the middle and is linear on both ends.
We can decompose Equation 13 into
$P\left(\left|\sum_{j=1}^{d}a_{j}\textup{median}(\tan(X_{j,1}),\cdots,\tan(X_{j,n}))-\sum_{j=1}^{d}a_{j}\right|>\epsilon\sum_{j=1}^{d}a_{j}\right)$
$=$ $P\left(\left|\sum_{j=1}^{d}a_{j}\tan(\textup{median}(X_{j,1},\cdots,X_{j,n}))-\sum_{j=1}^{d}a_{j}\right|>\epsilon\sum_{j=1}^{d}a_{j}\right)$
$\leq$ $P\left(\left|\sum_{j=1}^{d}a_{j}\tan(\textup{median}(X_{j,1},\cdots,X_{j,n}))-\sum_{j=1}^{d}a_{j}f(\textup{median}(X_{j,1},\cdots,X_{j,n}))\right|>0\right)$
$+$ $P\left(\left|\sum_{j=1}^{d}a_{j}f(\textup{median}(X_{j,1},\cdots,X_{j,n}))-\sum_{j=1}^{d}a_{j}\right|>\epsilon\sum_{j=1}^{d}a_{j}\right).$
For the first part, we have that
$P\left(\left|\sum_{j=1}^{d}a_{j}\tan(\textup{median}(X_{j,1},\cdots,X_{j,n}))-\sum_{j=1}^{d}a_{j}f(\textup{median}(X_{j,1},\cdots,X_{j,n}))\right|>0\right)$
$\leq$ $\sum_{j=1}^{d}P\left(\left|\tan(\textup{median}(X_{j,1},\cdots,X_{j,n}))-f(\textup{median}(X_{j,1},\cdots,X_{j,n}))\right|>0\right)$
$\leq$ $\sum_{j=1}^{d}P\left(\left|\textup{median}(X_{j,1},\cdots,X_{j,n})-\frac{\pi}{4}\right|>\Delta\frac{\pi}{4}\right)$
$\leq$ $2d\exp{(-\frac{n\Delta^{2}}{2})}$
where we apply Lemma D.1 in the last step.
For the second part, we first know that $\textup{median}(X_{j,1},\cdots,X_{j,n})-\frac{\pi}{4}$ is sub-Gaussian variable with $\sigma^{2}=\frac{\pi^{2}}{16n}$ from Lemma D.1. Then $f(\textup{median}(X_{j,1},\cdots,X_{j,n}))-\mathbb{E}f(\textup{median}(X_{j,1},\cdots,X_{j,n}))$ is sub-Gaussian variable with $\sigma^{2}=\frac{1}{\cos^{4}((1+\Delta)\frac{\pi}{4})}\frac{\pi^{2}}{16n}$ because $f^{\prime}(x)\leq\frac{1}{\cos^{2}((1+\Delta)\frac{\pi}{4})}$. And $\sum_{j=1}^{d}a_{j}f(\textup{median}(X_{j,1},\cdots,X_{j,n}))-\sum_{j=1}^{d}a_{j}\mathbb{E}f(\textup{median}(X_{j,1},\cdots,X_{j,n}))$ is sub-Gaussian variable with $\sigma^{2}=\frac{1}{\cos^{4}((1+\Delta)\frac{\pi}{4})}\frac{\pi^{2}}{16n}\left(\sum_{j=1}^{d}a_{j}\right)^{2}$.
When $n\geq\frac{\pi^{3}}{2\epsilon^{2}\cos^{4}((1+\Delta)\frac{\pi}{4})}$, it holds that
$\left|\mathbb{E}f(\textup{median}(X_{j,1},\cdots,X_{j,n}))-1\right|$ $\leq\mathbb{E}\left|f(\textup{median}(X_{j,1},\cdots,X_{j,n}))-f(\frac{\pi}{4})\right|$
$\leq\max_{x}f^{\prime}(x)\mathbb{E}\left|\textup{median}(X_{j,1},\cdots,X_{j,n})-\frac{\pi}{4}\right|$
$\leq\frac{1}{\cos^{2}((1+\Delta)\frac{\pi}{4})}\sqrt{2\pi}\frac{\pi}{4\sqrt{n}}\leq\frac{\epsilon}{2}$
and
$P\left(\left|\sum_{j=1}^{d}a_{j}f(\textup{median}(X_{j,1},\cdots,X_{j,n}))-\sum_{j=1}^{d}a_{j}\right|>\epsilon\sum_{j=1}^{d}a_{j}\right)$
$\leq$ $P\left(\left|\sum_{j=1}^{d}a_{j}f(\textup{median}(X_{j,1},\cdots,X_{j,n}))-\sum_{j=1}^{d}a_{j}\mathbb{E}f(\textup{median}(X_{j,1},\cdots,X_{j,n}))\right|>\frac{\epsilon}{2}\sum_{j=1}^{d}a_{j}\right)$
$+$ $\sum_{j=1}^{d}P\left(\left|\mathbb{E}f(\textup{median}(X_{j,1},\cdots,X_{j,n}))-1\right|>\frac{\epsilon}{2}\right)$
$\leq$ $2\exp{\left(-\frac{2n\epsilon^{2}\cos^{4}((1+\Delta)\frac{\pi}{4})}{\pi^{2}}\right)}.$
Combining these two parts, we can get that
$P\left(\left|\sum_{j=1}^{d}a_{j}\textup{median}(\tan(X_{j,1}),\cdots,\tan(X_{j,n}))-\sum_{j=1}^{d}a_{j}\right|\geq\epsilon\sum_{j=1}^{d}a_{j}\right)$
$\leq$ $2d\exp(-\frac{n\Delta^{2}}{2})+2\exp{\left(-\frac{2n\epsilon^{2}\cos^{4}((1+\Delta)\frac{\pi}{4})}{\pi^{2}}\right)}$
for $n\geq\frac{\pi^{3}}{2\epsilon^{2}\cos^{4}((1+\Delta)\frac{\pi}{4})}$. ∎